# A Comprehensive Survey on Deep Multimodal Learning with Missing Modality

### v2

@@ 2.1.1. Modality Composition Method
Figure 5. Description of two typical Modality generation methods. We set modality 2 as the missing modality for examples and use other available modalities to generate modality 2. “GEN” in both figures represents generative networks. (a) We set up a modality-2 generator (GEN-2) from other modalities. In the actual training process, modality 1 and 3 each have such a GEN. For simplicity, we omit the two GENs. (b) All modalities are input and generated together by using a single GEN.
2.1.2. Modality Generation Method
Various matrix imputation methods (Gong et al., 2021) were proposed in missing data research, leveraging latent correlations among matrix elements. However, missing data often appears in large blocks in multimodal datasets, making conventional methods inefficient for large-scale processing and high-dimensional computations. With deep learning, synthesizing missing modalities has become more efficient. Current methods for generating original data of missing modalities are divided into individual and unified generative methods. Individual modality generation method trains an individual generative model for each modality in case of any missing modality situation, which is shown in Figure 5(a). Early works used Gaussian processes (Mario Christoudias et al., 2010) or Boltzmann machines (Sohn et al., 2014) to generate missing modalities from available inputs. With deep learning, methods like AEs and U-Net(Ronneberger et al., 2015) can be used for original modality data generation. Li et al.(Li et al., 2014) used 3D-CNN to generate positron emission tomography (PET) data from magnetic resonance imaging (MRI) inputs. Chen et al.(Chen et al., 2024b) addressed missing modalities in MRI segmentation by training U-Net models to generate other two modalities from MRI. Recent work (Ma et al., 2021a) used AEs as one of the baselines to complete datasets by training one AE per modality. In domain adaptation, Zhang et al. (Zhang et al., 2021) proposed a Multi-Modality Data Generation module with domain adversarial learning to generate each missing modality by learning domain-invariant features. GANs significantly improved image generation quality by using a generator to create realistic data and a discriminator to distinguish it from real data. Researchers began replacing AEs and U-Nets with GANs for missing modality generation. For example, GANs generated original data of missing modalities using latent representations of existing ones in breast cancer prediction (Arya and Saha, 2021), and WGANs were applied in sentiment analysis (Wang et al., 2022). In remote sensing, Bischke et al.(Bischke et al., 2018) used GANs to generate depth data, improving segmentation over RGB-only models. GANs were also used in robotic recognition to generate RGB and depth images(Gunasekar et al., 2020). Recent studies (Ma et al., 2021a) show that GANs outperform AEs in generating more realistic missing modalities and can lead to better downstream-task model performance. Recently, the introduction of Diffusion models has further improved image generation quality. Wang et al. proposed the IMDer method (Wang et al., 2024a), which uses available modalities as conditions to help diffusion models generate missing modalities. Experiments showed diffusion reduces semantic ambiguity between recovered and missing modalities and achieves good generalization performance than previous works. But, training an individual generator for each modality is inefficient and time-consuming, failing to capture the latent correlations between modalities. Researchers develop another type of generative method, the unified modality generation method, which trains a unified model that can generate all modalities simultaneously (Figure 5(b)). One representative model is Cascade AE (Tran et al., 2017), which stacks AEs to capture the differences between missing and existing modalities for generating all missing modalities. Recent researchers, such as Zhang et al. (Zhang et al., 2023b), have attempted to use attention mechanisms and max-pooling layers to integrate features of existing modalities, enabling modality-specific decoders to generate each missing modality. Experiments demonstrated that this method is more effective than using max-pooling alone (Chartsias et al., 2017) to integrate features from any number of modalities for generating missing modalities. The above methods for generating original data for missing modalities can mitigate performance degradation to some extent. However, training a generator that can produce high-quality, real-world-distribution-like missing modalities remains challenging, especially when the training dataset contains a few full-modality samples. Additionally, the modality generation model significantly increases storage requirements. As the number of modalities grows, the complexity of these generative models also increases, further complicating the training process and resource demands.
Figure 6. Illustration of representation constraint methods (include regularization-based and correlation driven methods). Set modality 2 as the missing modality. Between the features of different modalities, correlation or regularization terms are used.
2.2. Feature Space Engineering
@@
The following introduces methods to address the missing modality problem at the feature space. First, we present two constraint-based approaches that enhance the learning of more discriminative and robust representations by imposing specific constraints (Figure 6). One method involves regularization to improve the effectiveness and generalization of the learned representations. Another method focuses on maximizing correlations, using specific measurements to strengthen the relationships between features. Next, the representation composition method, which can borrow the solutions discussed in Section 2.1.1, operates at the feature level of the modalities or employs arithmetic operations to handle a dynamic number of modalities. Finally, we introduce the representation generation method, which can generate the representations of missing modalities.
2.2.1. Regularization Based Method
There are few regularization based methods, such as Tensor Rank Regularization (Liang et al., 2019), which introduces tensor rank minimization as a regularization term in the training process. This method can automatically handle missing or noisy data without prior knowledge of data imperfections. It combines the temporal nonlinear transformation of multimodal data with simple regularization techniques of tensor structures, helping to learn true correlations and latent structures. This effectively mitigates the imperfection of input data and addresses the missing modality problem. Due to the limited number of this kind of methods, they have only been verified on time series data, and whether they can work in other fields requires extensive verification.
2.2.2. Correlation Driven Method
@@
Some methods train models by coordinating the correlation between different modal features. For example, Wang et al. (Wang et al., 2020a) use a Deep Canonical Correlation Analysis (CCA) module to maximize feature associations within the same category, enabling training on incomplete datasets. Ma et al. (Ma et al., 2021b) proposed a Maximum likelihood function to characterize conditional distributions of full-modality and missing-modality samples during training. Other researchers (Liu et al., 2021a) have added constraints based on the Hilbert-Schmidt Independence Criterion to help the model learn how to complete missing modality features from full modality features. These methods all aid models in learning how to complete missing modality features or training on incomplete datasets by learning the similarity or correlation between features. A drawback of above methods is that they perform well only when two or three modalities are used as input. Even when trained on a dataset with missing modalities, some methods still struggle to effectively address the missing modality problem during testing. Additionally, as the number of modalities increases, the variety of trade-offs in loss functions becomes more complex, making it challenging to achieve good generalization performance.
@@
Figure 7. The general idea of the arithmetic method. The representation of any number of modalities can be combined to a fixed dimension that is the same as the dimension required by the subsequent DNN layer.
2.2.3. Representation Composition Method
@@
There are two types of representation composition methods: one is the retrieval-based composition method, which attempts to recover the missing modality representation by borrowing existing samples or modality information, similar to the original data composition method in Section 2.1.1. They typically use pre-trained feature extractors to generate features from available samples, storing them in a feature pool (Zhi et al., 2024; Sun et al., 2024a). Cosine similarity is then used to retrieve matched features for the input sample, with the average of these missing-modality features in top-k samples filling the representation of missing modalities. Additionally, some methods, such as Missing Modality Feature Generation (Wang et al., 2023c), replace missing modality features by averaging the representations of available modalities, assuming similar feature distributions across modalities. Another representation composition method, the arithmetic method, dynamically processes modality representations without using learnable parameters like pooling layers (Figure 7). Researchers (Zhou et al., 2021; Delgado-Escano et al., 2021; Marín-Jiménez et al., 2021; Wang et al., 2023c) have fused features through operations like average/max pooling or addition, offering low computational complexity and efficiency. However, drawbacks include potential loss or dilution of important information. To address this, merge operations (Delgado-Escano et al., 2021; Marín-Jiménez et al., 2021) use a sign-max function, which selects the highest absolute value from feature vectors, yielding better results by preserving both positive and negative activation values. TMFormer (Zhang et al., 2024) introduces token merging based on cosine similarity, while others calculate moments of available modalities for feature fusion through addition and concatenation (Havaei et al., 2016). Similar approaches (Zheng et al., 2024a, b; Zhang et al., 2024) focus on selecting key vectors or merging tokens to handle missing modalities. Recent works (Zheng et al., 2024a, b) focus on selecting key features or vectors for handling missing modalities. Despite these advancements, methods can still struggle if a major modality is missing. Therefore, those representation composition approaches not only allow the model to flexibly handle features from any number of modalities but also enable learning without/few learnable parameters. It is also difficult to capture the inter-modality relationships through the learning way, as methods like selecting the largest vector or the feature with the highest score is difficult to fully represent the characteristics of all modalities.
(a) Indirect-to-Task Representation Generation Method
(b) Direct-to-Task Representation Generation Method
@@
Figure 8. Description of two typical representation generation methods. We assume modality 2 is missing here. (a) This kind of methods are supervised by the loss of two tasks, auxiliary reconstruction and downstream tasks. Note that these two tasks sometimes can be trained separately, first training the reconstruction task, discarding the generator, and then training the downstream task, such as the training progress of ActionMAE (Woo et al., 2023). While (b) methods directly use a generative network (“GEN-2” in figure) to generate representations of the missing modality 2 during training and inference without reconstructing the original data of modality 2.
2.2.4. Representation Generation Method
@@
Compared to original data generation methods in Section 2.1.2, hierarchical representation generation can integrate seamlessly with existing multimodal frameworks. Current methods fall into two categories: (1) Indirect-to-task representation generation methods (Figure 8(a)) treat modality reconstruction as an auxiliary task during training, helping the model intrinsically generate missing modality representations for downstream tasks. Since the auxiliary task aids in representation generation during training but is dropped during inference of downstream task, it is termed “indirect-to-task.” (2) Direct-to-task representation generation methods (Figure 8(b)) train a small generative model to directly map available modality information to the representation of missing modalities.
@@
Indirect-to-Task Representation Generation Methods: A common architecture for indirectly generating missing modality representations is the “encoder-decoder” model. During training, modality-specific encoders extract features from available modalities, one/some reconstruction decoders reconstructs the missing modality, and one downstream module supervised by downstream task loss for prediction. Those reconstruction decoders are discarded during inference, where predictions rely on both the generated missing modality representations and the existing ones. This approach typically employs Multi-Task Learning, training both downstream tasks and auxiliary tasks (modality reconstruction) simultaneously. Some methods, inspired by Masked Autoencoder (He et al., 2022), split reconstruction and downstream task training. Based on the input features used by reconstruction decoders, we divide these methods into two categories. The first category, Before Fusion methods, inputs the features from modality-specific encoders directly into each specific reconstruction decoder for missing modality reconstruction. For example, MGP-VAE (Hamghalam et al., 2021) uses the VAE encoder’s output for Gaussian Process prediction and reconstruction decoders. Similarly, PFNet (Zheng et al., 2021) employs RGB and thermal-infrared modalities for pedestrian re-identification, feeding outputs of each encoder into corresponding decoders before fusing them together. Li et al. (Li et al., 2024b) adopted a similar approach for brain tumor segmentation, reconstructing each modality before fusing encoder outputs. The second category, After Fusion methods, uses the fused features from modality-specific encoders as the input of reconstruction decoder. Tsai et al.(Tsai et al., 2018) introduced a Multimodal Factorization Model where modality-specific encoder-decoders reconstruct the missing modality, while the separate downstream task model employs a multimodal encoder-decoder. To recover the the missing modality features, the outputs of the downstream task model’s encoder are not only fused with the outputs of each modality-specific encoder for inputting to the each corresponding reconstruction decoder, but also used for predictions of the downstream task decoder. Chen et al.(Chen et al., 2019) also decoupled the encoders for reconstruction and downstream tasks, showing that disentangling features helped eliminate modality-specific information. Jeong et al.(Jeong et al., 2022) proposed a simpler approach using multi-level skip connections in a U-Net architecture to fuse features for missing modality reconstruction, achieving better generalization in brain tumor segmentation than Chen et al. Further, Zhou et al.(Zhou, 2023) and Sun et al. (Sun et al., 2024b) demonstrated effective cross-modal fusion techniques to enhance reconstruction and segmentation. Most of these methods train reconstruction and downstream tasks concurrently. However, some approaches, such as ActionMAE (Woo et al., 2023) and M3AE (Liu et al., 2023c), first train reconstruction tasks and then fine-tune the pretrained encoders for downstream tasks. Despite achieving strong generalization, these methods are limited by the requirement of supervising by the original data of missing modalities.
@@
Direct-to-Task Representation Generation Methods: In contrast to indirect methods, these approaches directly map information from existing modalities to the missing modality representation. The general concept is illustrated in Figure 8(b). Early work (Mario Christoudias et al., 2010) using Gaussian processes to “hallucinate” missing modalities inspired Hoffman et al. (Hoffman et al., 2016) to propose a Hallucination Network (HN) that predicts missing depth modality features from RGB images. The HN (student) aligns intermediate layer features with a teacher network (trained on depth images) using an L2 loss, allowing RGB inputs to generate depth features to address missing modality challenges. This method has been extended to pose estimation (Choi et al., 2017) and land cover classification (Kampffmeyer et al., 2018), where the HN generates heat distribution and depth features. In tasks like sentiment analysis, a translation module can map representations of available modalities to the missing ones, with L2 loss applied to align outputs with those from networks trained on missing modalities (Huan et al., 2023). A notable recent work, SMIL (Ma et al., 2021a), employs Bayesian learning and meta-regularization to directly generate missing modality representations. SMIL has shown strong generalization on datasets with a high rate of missing modality samples, handling up to 90% missing-modality samples. Subsequent work (Li et al., 2022c) addressed challenges such as distribution inconsistency, failure to capture specific modality information, and lack of correspondence between generated and existing modalities. They introduced solutions such as discriminator-based consistency loss, memory representativeness, and structural alignment. Researchers have also explored cross-modal distribution transformation methods to align distributions of generated and existing modality representations, enhancing discriminative ability (Wang et al., 2023b). In audio-visual question answering (AVQA), Park et al. (Park et al., 2024) proposed a Relation-aware Missing Modal Generator to generate pseudo features of missing modalities, improving robustness and accuracy. Compared to indirect methods, direct generation avoids the cumbersome reconstruction of raw data, using a feature generator that is easier to train and integrate into multimodal models. Recently, Kim et al.(Kim and Kim, 2024) combined self-supervised learning, Joint Embedding Predictive Architecture, and parameter-efficient learning to help the model generate missing modality features. Additionally, Guo et al.(Guo et al., 2024) proposed a Missing Modality Generation Module for prompt learning, mapping prompts of available modalities to those of missing ones. Lin et al. (Lin et al., 2024) introduced U-Adapter, which performs cross-modal fusion while preventing unreliable information due to domain shifts, effectively solving the missing modality problem.
@@
In summary, above representation generation methods are all limited by the imbalance of missing modalities in the training dataset or the limited number of full-modality samples, which may lead to the possibility of being dominated by existing modalities. In addition, since indirect-to-task representation generation methods can access the missing modality data of the missing modality samples during training, which means that the training dataset is actually complete, this type of method usually performs better than direct-to-task representation generation methods.
2.3. Architecture Engineering

### v3

@@ 3.1.1. Modality Composition Methods
Figure 5. Description of two typical modality generation methods. We set modality 2 as the missing modality for examples and use other available modalities to generate modality 2. “GEN” in both figures represents modality generation networks. (a) We set up a modality-2 generator (GEN-2) from other modalities. (b) All modalities are input and generated together by using a single GEN.
3.1.2. Modality Generation Methods
With deep learning, synthesizing missing modalities has become more effective by leveraging powerful representation learning and generative models that can capture complex cross-modal relationships. Current methods for generating missing-modality data are divided into individual and unified generative methods.
Individual Modality Generation Methods train an individual generative model for each modality in case any modality is missing, as shown in Figure 5(a). Early works used Gaussian processes (Mario Christoudias et al., 2010) or Boltzmann machines (Sohn et al., 2014) to generate missing modalities from available data. With deep learning, models like Auto-Encoders (AEs) (Hinton and Zemel, 1993) can be used for missing modality generation. Li et al. (Li et al., 2014) used 3D-CNN to generate positron emission tomography (PET) data from magnetic resonance imaging (MRI) inputs. Chen et al. (Chen et al., 2024b) addressed missing modalities in MRI segmentation by training U-Net models to generate other two modalities from MRI. A recent work (Ma et al., 2021b) used AEs as one of the baselines to complete datasets by training one AE per modality. In domain adaptation, Zhang et al. (Zhang et al., 2021) proposed a U-Net-based Multi-Modality Data Generation module with domain adversarial learning to generate each missing modality by learning domain-invariant features.
GANs significantly improved image generation quality by using a generator to create realistic data and a discriminator to distinguish it from real data. Researchers therefore began replacing above models with GANs for modality generation. For example, GANs generate missing modalities using latent representations of existing ones in breast cancer prediction (Arya and Saha, 2021). In remote sensing, Bischke et al. (Bischke et al., 2018) used GANs to generate depth data, improving segmentation over RGB-only models. GANs were also used in robotic object recognition models to help generate missing RGB and depth images(Gunasekar et al., 2020). Recent studies (Ma et al., 2021b) show that GANs outperform AEs in generating more realistic missing modalities and can lead to better downstream-task model performance. Recently, the introduction of Diffusion models has further improved image generation quality. Wang et al. proposed the IMDer method (Wang et al., 2024a), which uses available modalities as conditions to make diffusion models generate missing modalities. Experiments showed it can reduce semantic ambiguity between recovered and missing modalities and achieves good generalization performance than previous works.
Unified Original Data Generative Methods train a unified model that can generate all modalities simultaneously (Figure 5(b)). One representative model is Cascade AE (Tran et al., 2017), which stacks AEs to capture the differences between missing and existing modalities for generating all missing modalities. Recently, Zhang et al. (Zhang et al., 2023b), have attempted to use attention and max-pooling to integrate features of existing modalities, enabling modality-specific decoders to accept available modalities and generate other missing modalities. Experiments demonstrated that this method is more effective than using max-pooling alone (Chartsias et al., 2017) to integrate features from any number of modalities for generating all missing modalities together. Above methods for generating missing modalities can mitigate performance degradation to some extent. However, when facing a scenraio where one of the modalities is severely missing, training a generator that can produce high-quality missing modalities remains challenging. Additionally, the modality generation model significantly increases storage and computational requirements. As the number of modalities grows, the complexity of these generative models also increases, further complicating the training process and resource demands.
Figure 6. Illustration of coordinated-representation methods, which impose constraints to make the learned representations semantically consistent. Assuming that modality 2 is missing, these methods use constraints terms between the features of different modalities.
3.2. Representation-Focused Models
@@
Representation-focused models address the missing modality problem at the representation level. We introduce such models by first presenting two coordinated-representation-based approaches that enhance the learning of more discriminative and robust representations by imposing specific constraints (Figure 6). The next type of representation-focused methods that we discuss are the representation imputation methods, which can be categorized into representation composition and representation generation methods. Representation composition methods can borrow the solutions described in Section 3.1.1 and operate at the representation level of the modalities or employs arithmetic operations (e.g., pooling) to fuse a dynamic number of modalities. Finally, we introduce the representation generation methods, which usually use small generative models to produce the representations of missing modalities.
3.2.1. Coordinated-Representation Methods
Coordinated-representation methods focus on introducing certain constraints between representations of different modalities to make the learned representations semantically consistent. They are usually divided into two categories, one based on regularization and the other driven by correlation. An example of regularization is Tensor Rank Regularization (TRR) (Liang et al., 2019). TRR achieves multimodal fusion by taking the outer products of different-modality tensors and taking the sum of all their outer products. In order to solve the high rank of multimodal representation tensors caused by imperfect modalities (missing or noisy) in time series, Paul et al. introduced tensor rank minimization to try to keep the multimodal representation low rank to better express the tensors of true correlation and potential structure in multimodal data, thereby alleviating the imperfection of the input.
@@ 3.2.1. Coordinated-Representation Methods
Some methods train models by coordinating the correlation between different modal features. For example, Wang et al. (Wang et al., 2020a) use a Deep Canonical Correlation Analysis (CCA) module to maximize feature associations of available modalities by using Canonical Correlation Coefficient, which enables training on incomplete datasets. Ma et al. (Ma et al., 2021c) proposed a Maximum likelihood function to characterize conditional distributions of full-modality and missing-modality samples during training. Other research efforts (Liu et al., 2021a) have added constraints based on the Hilbert-Schmidt Independence Criterion (HSIC), which helps models learn how to complete missing modality features by enforcing independence between irrelevant features while maximizing the dependence between relevant ones. These methods (Matsuura et al., 2018; Ma et al., 2021a; Li et al., 2022a) aid models in learning how to complete missing modality features or training on incomplete datasets by learning the similarity or correlation between features. A drawback of above methods is that they perform well only when two or three modalities are used as input (Zhao et al., 2024). Even when trained on a dataset with missing modalities, some methods still struggle to effectively address the missing modality problem during testing.
@@
Figure 7. The general idea of the arithmetic operation-based representation composition methods. The representation of any number of modalities can be combined to a fixed dimension that is the same as the dimension required by the subsequent DNN layer.
3.2.2. Representation Composition Methods
There are two types of representation composition methods, which are explained below. Retrieval-Based Representation Composition Methods attempt to recover the missing modality representation by retrieving the modality data from existing samples, similarly to the modality composition method in Section 3.1.1. They typically use pre-trained feature extractors to generate features from available samples, storing them in a feature pool (Zhi et al., 2024; Sun et al., 2024a). Cosine similarity is then used to retrieve matched features for the input sample, with the average of these missing-modality features in top-K samples filling the representation of missing modalities. Additionally, some methods, such as Missing Modality Feature Generation (Wang et al., 2023c), replace missing modality features by averaging the representations of available modalities, assuming similar feature distributions across modalities.
@@ 3.2.2. Representation Composition Methods
Arithmetic Operation-Based Representation Composition Methods can flexibly fuse any number of modality representations through arithmetic operations such as pooling without learnable parameters (Figure 7). Researchers (Zhou et al., 2021; Wang et al., 2023c) have fused features through operations like average/max pooling or addition, offering low computational complexity and efficiency. However, drawbacks include potential loss of important information. To address this, merge operations (Delgado-Escano et al., 2021) use a sign-max function, which selects the highest absolute value from feature vectors, yielding better results by preserving both positive and negative activation values. TMFormer (Zhang et al., 2024) introduces token merging based on cosine similarity. Similar approaches (Zheng et al., 2024) focus on selecting key vectors or merging tokens to handle missing modalities.
Therefore, those representation composition approaches not only allow the model to flexibly handle features from any number of modalities but also enable learning without introducing new learnable parameters. However, they are not good at capturing the inter-modality relationships through learning, as methods like selecting the largest vector or the feature with the highest score is difficult to fully represent the characteristics of all modalities.
(a) Indirect-to-Task Representation Generation Methods
(b) Direct-to-Task Representation Generation Methods
@@
Figure 8. Description of two typical representation generation methods. We assume modality 2 is missing here. (a) Indirect-to-task representation generation methods are supervised by the loss of two tasks, namely: auxiliary reconstruction and downstream tasks. Note that these two tasks sometimes can be trained separately, first training the reconstruction task, discarding the generator, and then training the downstream task, such as the training progress of ActionMAE (Woo et al., 2023). While (b) direct-to-task representation generation methods directly use a modality generation network (“GEN-2” in figure) to generate representations of the missing modality 2 during training and inference without reconstructing the modality 2.
3.2.3. Representation Generation Methods
@@
Compared to modality generation methods in Section 3.1.2, representation generation methods can be integrated seamlessly into existing multimodal frameworks. Current methods fall into two categories: (1) Indirect-to-task representation generation methods (Figure 8(a)) treat modality reconstruction as an auxiliary task during training, helping the model intrinsically generate missing modality representations for downstream tasks. Since the auxiliary task aids in representation generation during training but is dropped during inference of downstream task, it is termed “indirect-to-task.” (2) Direct-to-task representation generation methods (Figure 8(b)) train a small generative model to directly map available modality representations to the missing modalities. We provide more details about these two categories of methods below.
Indirect-to-Task Representation Generation Methods can indirectly generate missing modality representations and often employ the “encoder-decoder” architecture. During training, modality-specific encoders extract features from available modalities, reconstruction decoders reconstruct the missing modality, and the downstream module is supervised by downstream task loss for prediction. Those reconstruction decoders are discarded during inference, where predictions rely on both the generated missing modality representations and the existing ones. This approach typically employs Multi-Task Learning, training both downstream tasks and auxiliary tasks (modality reconstruction) simultaneously. Some methods, inspired by Masked Autoencoder (He et al., 2022), split reconstruction and downstream task training. Typically, in the training process of this type of methods, the features of the available modalities are usually input into the downstream task module and the reconstruction decoder of different modalities after being fused. However, some methods accept the outputs of different modality encoders into their respective reconstruction decoders. Based on this distinction, we divide this type of methods into before-fusion and after-fusion methods.
Before Fusion methods receive the features from modality-specific encoders directly into each specific reconstruction decoder for missing modality reconstruction. For example, MGP-VAE (Hamghalam et al., 2021) uses output of the VAE encoder of each modality as the input of each reconstruction decoder and the fused output for Gaussian Process prediction. Similarly, PFNet (Zheng et al., 2021) employs RGB and thermal-infrared modalities for pedestrian re-identification, feeding outputs of each encoder into corresponding decoders before fusing them together. Li et al. (Li et al., 2024b) adopted a similar approach for brain tumor segmentation, reconstructing each modality before fusing encoder outputs.
@@
After Fusion methods use the fused features from modality-specific encoders as the input of reconstruction decoder. Tsai et al. (Tsai et al., 2018) introduced a Multimodal Factorization Model where modality-specific encoder-decoders reconstruct the missing modality, while the downstream task module employs a separate multimodal encoder-decoder. To recover the missing modality features, the outputs of the downstream task model’s encoder are not only used for predictions of the downstream task decoder, but also fused with the outputs of each modality-specific encoder for inputting to the each corresponding reconstruction decoder. Chen et al. (Chen et al., 2019) also decoupled the encoders for reconstruction and downstream tasks, showing that disentangling features helped eliminate irrelevant features. Jeong et al. (Jeong et al., 2022) proposed a simpler approach using multi-level skip connections in a U-Net architecture to fuse features for missing modality reconstruction, achieving better generalization in brain tumor segmentation than Chen et al. (Chen et al., 2019). Further, Zhou et al. (Zhou, 2023) and Sun et al. (Sun et al., 2024b) demonstrated effective cross-modal attention fusion techniques could help enhance reconstruction and segmentation. Most of these methods train reconstruction and downstream tasks concurrently. However, some approaches, such as ActionMAE (Woo et al., 2023) and M3AE (Liu et al., 2023c), first train reconstruction tasks and then fine-tune the pre-trained encoders for downstream tasks. Although these methods achieve strong performance, they still require the missing modality data to be present during training to supervise the reconstruction decoders.
Direct-to-Task Representation Generation Methods usually use simple generative models to directly map information from avaiable modalities to the missing modality representation. The general concept is illustrated in Figure 8(b). Early work (Mario Christoudias et al., 2010) using Gaussian Processes to “hallucinate” missing modalities inspired Hoffman et al. (Hoffman et al., 2016) to propose a Hallucination Network (HN) that predicts missing depth modality features from RGB images. The HN aligns intermediate layer features with another network (trained on depth images) using an L2 loss, allowing RGB inputs to generate depth features to address missing modality challenges. This method has been extended to pose estimation (Choi et al., 2017) and land cover classification (Kampffmeyer et al., 2018), where the HN can be used to generate missing heat distribution and depth features.
@@
In tasks like sentiment analysis, a translation module can map available-modality representations to the missing ones, with L2 loss applied to align outputs with those from networks trained on missing modalities (Huan et al., 2023). A notable recent work, SMIL (Ma et al., 2021b), employs Bayesian learning and meta-regularization to directly generate missing modality representations. SMIL has shown strong generalization on datasets with a high rate of missing modality samples. Subsequent work (Li et al., 2022d) addressed challenges such as distribution inconsistency, failure to capture specific modality information, and lack of correspondence between generated and existing modalities. Researchers have also explored cross-modal distribution transformation methods to align distributions of generated and existing modality representations, enhancing discriminative ability (Wang et al., 2023b). In audio-visual question answering (AVQA), Park et al. (Park et al., 2024) proposed a Relation-aware Missing Modal Generator (RMM) that consist of to generate pseudo features of missing modalities, improving robustness and accuracy. The RMM consists of a visual generator, auditory generator, and text generator, which generate missing auditory representations by analyzing the correlations between visual and textual modalities and utilizing learnable parameter vectors (slots) for reconstruction and synthesis of missing-modality features.
Recently, to address the potential absence of different modalities in sentiment analysis, Guo et al. (Guo et al., 2024) proposed a Missing Modality Generation Module for prompt learning. This module maps prompts of available modalities to prompts of missing modalities through a set of projection layers. Lin et al. (Lin et al., 2024) proposed Uncertainty Estimation Module which can identify useless tokens in different modality tokens to facilitate better use of U-Adapter-assisted pre-trained models to better utilize the tokens of available modalities when downstream tasks face missing modalities.
Compared to indirect methods, direct generation methods avoid the cumbersome reconstruction of modality data, using feature generators that are easier to train and integrate into multimodal models.
@@
In summary, representation generation methods are limited by the imbalance of missing modalities in the training dataset or the limited number of full-modality samples, which may lead to training procedures that overfit to the existing modalities. In addition, indirect-to-task representation generation methods can access the missing modality data of the missing modality samples during training, which means that the training dataset is actually complete, so this type of method usually performs better than direct-to-task representation generation methods.
4. Methodologies in Strategy Design Aspect
4.1. Architecture-Focused Models